OpenStudy (anonymous):

Write in Logarithmic Form: 125 = 5^3

OpenStudy (cwrw238):

if x = a^b then b = log a x

OpenStudy (cwrw238):

or if you prefer numbers 100 = 10 ^ 2 2 = log10 100 10 is called the base of the log and the log itself is the power to which the base is taken

OpenStudy (cwrw238):

hope this helps

OpenStudy (anonymous):

thank you!

OpenStudy (cwrw238):

yw